this is a word problem
20 less than twice a number is equal to the number. What is the number?

Answer by checkley71(8403)

-6/19 IS A RATIONAL NUMBER
AN IRRATIONAL NUMBER IS FOLLOWED BY i.
-----------------------------------------------
2X-20=X
2X-X=20
X=20
PROOF
2*20-20=20
40-20=20
20=20